Calculate the acceleration of Josh riding his bicycle in a straight line that speeds up from

4 m/s to 6 m/s in 5 seconds.


1
Expert's answer
2021-03-14T19:12:35-0400
"a=\dfrac{v-v_0}{t}=\dfrac{6\ \dfrac{m}{s}-4\ \dfrac{m}{s}}{5\ s}=0.4\ \dfrac{m}{s^2}."